# Electric-Field Driven Nuclear Dynamics of Liquids and Solids from a Multi-Valued Machine-Learned Dipolar Model

### v1

@@ Abstract
The driving of vibrational motion by external electric fields is a topic of continued interest, due to the possibility of assessing new or metastable material phases with desirable properties. Here, we combine ab initio molecular dynamics within the electric-dipole approximation with machine-learning neural networks (NNs) to develop a general, efficient and accurate method to perform electric-field-driven nuclear dynamics for molecules, solids, and liquids. We train equivariant and autodifferentiable NNs for the interatomic potential and the dipole, modifying the prediction target to account for the multi-valued nature of the latter in periodic systems. We showcase the method by addressing property modifications induced by electric field interactions in a polar liquid and a polar solid from nanosecond-long molecular dynamics simulations with quantum-mechanical accuracy. For liquid water, we present a calculation of the dielectric function in the GHz to THz range and the electrofreezing transition, showing that nuclear quantum effects enhance this phenomenon. For the ferroelectric perovskite LiNbO3, we simulate the ferroelectric to paraelectric phase transition and the non-equilibrium dynamics of driven phonon modes related to the polarization switching mechanisms, showing that a full polarization switch is not achieved in the simulations.
@@ I Introduction
The interaction of matter with static and dynamical external electric fields plays a fundamental role in understanding the behavior of atoms, molecules, and complex materials. This interaction is exploited in fields as diverse as semiconductor nanotechnology [1] and enzymatic reactions [2]. In particular, there is a rising interest in tuning static and low-frequency dielectric fields to drive reversible phase transitions of materials, with the goal of controlling mechanical properties, increasing the efficiency of ionic conductors and capacitors, or making better energy or memory storage devices [1, 2, 3]. The large spatial extent and, depending on the phenomenon, the time-scale of many nanoseconds involved in the nuclear dynamics that define the response of a material to the application of such electric fields has been a long-standing challenge for quantum-mechanical first-principles atomistic simulations. While the physics of the matter-field coupling is known and appropriate simulation techniques already exist within the context of quantum-mechanical theories [4, 5, 6], simplified and non-transferable models [3] have often been used for describing these quantum phenomena at larger system sizes and longer time scales in complex systems.
@@
The goal of this work is to achieve a general framework that: (i) is generally applicable to different material classes (e.g. molecules, solids, liquids, and disordered systems) without relying on system-specific assumptions; (ii) keeps quantum-mechanical accuracy for electronic interactions and can include nuclear quantum effects; (iii) allows performing nanoseconds-long simulations including thousands of atoms in and out of equilibrium. It is clear that the rapidly evolving techniques of machine-learning (ML) in atomistic simulations can be applied to this problem. ML models have already been proven to succeed in the prediction of energies, forces and electronic-structure properties of materials [7, 8] with an accuracy comparable to that of the underlying quantum electronic-structure method, but at a small fraction of the cost and often with a strongly reduced scaling with system size. A fundamental quantity to describe the interaction of matter with electric fields is the dipole and its derivative with respect to nuclear displacements. Several works have already proposed ML-based models for these quantities [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 13]. While these approaches are appropriate and successful in the investigated examples, i.e., close to structural equilibria, they can become inaccurate when large displacements take place. As we will show in this work, an additional term arising from the multi-valued nature of the dipole in periodic systems sensitively influences the applicability of the ML model in these cases.
@@
By explicitly incorporating the multi-valued nature of the dipole in our model, we achieve a framework that is applicable to neutral and polar materials at different thermodynamic state points and far from equilibrium. We showcase the capabilities of our model by addressing liquid water and the solid ferroelectric perovskite LiNbO3 [20]. We calculate nanoseconds of first-principles-quality molecular dynamics coupled to static and time-dependent electric fields for these polar materials. For water, we show simulations of the dielectric function [21] spanning the GHz to the THz range and of the electrofreezing transition [22, 23], finding an enhancement of this phenomenon due to nuclear quantum effects at larger field intensities. For LiNbO3, we show the calculation of the ferroelectric to paraelectric phase transition [24, 25, 26, 27] and of the full-dimensional non-equilibrium and non-linear driving of phonon-modes in this material. For this last example, incorporating the multi-valued nature of the dipole in the ML model is essential. It shows that full-dimensional simulations can better explain the experimental observation that an ultrafast polarization switch is only transiently and partially achieved [28, 29, 30].
@@
In the following, we will summarize the theory of the model we propose, explain the ML model we develop, and discuss the rich phenomenology that these simulations can address together with the novel physical insights they can bring.
II Theory of Molecular Dynamics Driven by Electric Fields
When the applied electric field $\bm{\mathcal{E}}$ is small compared to the internal fields, one can apply the electric dipole approximation (EDA) within the linear-response regime (see Supplementary Section S1). In this case, the nuclear Hamiltonian can be expressed as [4, 6, 5]
[…]
where $T(\bm{p})$ is the kinetic energy, $\bm{p}$ the nuclear momenta, $\bm{R}$ the nuclear coordinates, $V(\bm{R})\equiv E[n_{0}]$ is the Born-Oppenheimer potential energy surface (PES) obtained from the ground-state electronic density $n_{0}$, and $\bm{\mu}(\bm{R})\equiv\bm{\mu}[n_{0}]$ is the system dipole.
The forces on the nuclei depend on the external field $\bm{\mathcal{E}}$ through
[…]
The first term on the right-hand side is the usual Born-Oppenheimer (BO) force. The second term can be expressed in terms of the atomic-polar tensors [17, 31], also called Born effective charges (BEC) $\bm{Z}^{*}$,
$\bm{Z}^{*}=\,\frac{1}{e}\frac{\partial\bm{\mu}}{\partial\bm{R}}=\frac{1}{e}\frac{\partial\bm{F}}{\partial\bm{\mathcal{E}}}$ (3)
[…]
For a given atomic configuration, the BEC can be computed from any of the two expressions in Eq. (3). The (relative) efficiency of these approaches strongly depends on the system, on the chosen electronic-structure code, and on details of the implementations. Evidently, the expressions can be calculated by finite differences or by perturbation theory [32, 33]. While the latter is elegant and can be quite efficient especially for systems where symmetry can be exploited, the former ensures a natural handling of electronic-structure level of approximation, such as different density-functionals, relativistic corrections and other terms that are cumbersome to evaluate analytically. In any event, the calculation of $\bm{Z}^{*}$ always depends on the calculation of $\bm{\mu}$, and thus represents an added cost on the already costly electronic-structure calculation.
We aim to circumvent the expensive first-principles calculation of $\bm{Z}^{*}$ that enter Eq. (2) by developing a ML model that is trained only on $\bm{\mu}$ and that is autodifferentiable [34] with respect to its input quantities, the nuclear coordinates $\bm{R}$. Training on $\bm{\mu}$ presents some advantages, including the ease to produce data for this quantity with diverse electronic-structure methods and the guarantee that the acoustic sum rule [35] for $\bm{Z}^{*}$ in the model is preserved, making it strictly translationally invariant.
II.1 Dipoles in Aperiodic and Periodic Systems
[…]
In periodic, condensed-phase systems, $\bm{\mu}_{e}$ cannot be directly evaluated by the expression above, as restricting the integration domain to the primitive unit cell would yield boundary-sensitive results. The dipole is related to the polarization $\bm{P}$ of a periodic system, in that the polarization is the dipole per unit volume, $\bm{P}=\bm{\mu}/\Omega$. The modern theory of polarization (MTP) [36, 37, 38] provides a proper definition of this quantity in periodic systems, showing that the electronic contributions to $\bm{P}$ come from many-body effects associated to the geometric phase of the wavefunction. There are several different (but equivalent) ways to evaluate $\bm{\mu}_{e}$ within MTP, which are all related to a way of evaluating the Berry connection [37, 38, 39]. As a consequence, the components of $\bm{\mu}$ corresponding to each lattice vector $\bm{a}_{\alpha}$ can be defined only up to quantum of polarization $e\left|\bm{a}_{\alpha}\right|$ and the dipole along $\bm{a}_{\alpha}$ is
$\mu_{\alpha}\equiv\mu_{\alpha}\mod e|\bm{a}_{\alpha}|.$ (5)
@@ II.1 Dipoles in Aperiodic and Periodic Systems
Because of the congruence relation in Eq. (5), the dipole is better represented as a lattice of vectors [38], and only differences of $\mu_{\alpha}$ values can be well defined. If one imposes translational invariance on the dipole by enforcing $\bm{\mu}$ to assume the same value at translationally-equivalent periodic atomic sites, the function $\bm{\mu}(\bm{R})$ becomes discontinuous because this implies “switching branches” across the periodic boundaries, as shown in Fig. 1a. Instead, a smooth and continuous function is obtained by continuing to follow the same branch, (Fig. 1a). Since the dipole then assumes different values for translationally-equivalent atomic positions, $\bm{\mu}(\bm{R})$ does not have translation invariance with respect to the periodic boundary conditions.
@@
In other words, a consequence of the multi-valued nature of $\bm{\mu}$ is that closed integrals do not vanish, but can take up (multiple) polarization quanta, depending on how often the periodic boundaries are crossed. We can quantify the change of the dipole $\Delta\bm{\mu}$ along a closed path $\gamma$ by a line integral. The result of this operation is non-trivial [40]:
@@
we can describe the more general case case where $N_{\rm at}$ atoms of the same oxidation state are displaced along the same direction. $\mathcal{N}_{I}$ is the difference between the values of $\mathcal{N}(s)$ upon crossing the periodic boundary. The values obtained for $\mathcal{N}_{I}$ are reported in Fig. 1b and c, and follow the expected oxidation numbers in these systems.
II.2 Machine-learning model for dipoles
The previous considerations lead to practical consequences when building a ML model for $\bm{\mu}(\bm{R})$ of a periodic system. Currently, equivariant neural networks (ENNs) yield the best accuracy to machine-learn properties of materials [43]. Most commonly, ENNs take as inputs the tuple $\left(Z;\bm{R},\bm{a}_{\alpha}\right)$, where $Z$ is the atomic number, and model the target quantity as a function of the interatomic distances $\Delta\bm{R}_{IJ}$, which are evaluated within the minimum image convention. We will call an atomic environment $\bm{\mathcal{A}}$ an element of the configuration space, which is isomorphic to a $3N$-dimensional torus $\mathbb{T}^{3N}$.
Going back to Fig. 1a, because such models require that the function to be learned is differentiable, the discontinuous curve which is obtained by matching $\bm{\mu}$ to assume the same value at equivalent $\bm{\mathcal{A}}$ cannot be learned in its full domain. If one is only interested in small atomic displacements close to equilibrium, a usual ML model will provide accurate results for $\bm{\mu}$ in this restricted space. However, as one moves into $\bm{\mathcal{A}}(\bm{R})$ that are far from equilibrium, the ML model will start to change its slope, in order to obey $\bm{\mu}(\bm{\mathcal{A}}(\bm{R}+\bm{l}))=\bm{\mu}(\bm{\mathcal{A}}(\bm{R}))$. In an autodifferentiable model, this results in $\bm{Z}^{*}$ and consequently forces that start to deviate from the correct value. If more data including environments far from equilibrium are added to the training set, the model will be unable to learn (see Supplementary Sections S2 for a practical example).
@@ II.2 Machine-learning model for dipoles
where $I$ runs over all atoms of the system and $\tilde{\bm{\mu}}$ can be handled by the usual learning procedure, as it depends only on $\bm{\mathcal{A}}$. From the definition of $\tilde{\bm{\mu}}^{\text{MV}}$ and $\bm{Z}^{*}$ in Eqs. (3) and (9), we arrive at
@@
In the implementation we used in this work, we modified the output of the forward pass of MACE to include the second term on the right of Eq. (9). In this way, we always calculate $\tilde{\bm{\mu}}^{\text{MV}}$ in the training and use that also in the calculation of the loss-function. Training is successful and accurate even for datasets containing very out of equilibrium structures of H2O and LiNbO3, and $\bm{Z}^{*}$ obtained by autodifferentiation is also accurate (see Supplementary Section S2 and Section V).
@@
The successes of previous ML models of dipoles and BEC are numerous [10, 14, 11, 46, 47, 18, 48, 16, 19, 9, 17]. They are successful because they either treat diffusive systems where the combined $\mathcal{N}$ (summing over all diffusing atoms) is zero, or address problems where the displacements along “ferroelectric” modes are not large and the system stays close to equilibrium along that coordinate. We note that models targeting to learn $\bm{Z}^{*}$ can also be problematic far from equilibrium if they miss the last term in Eq. (10) and they are not guaranteed to obey the acoustic sum rule for this quantity. The approach we present encompasses and extends these models to be generally applicable. This is for instance important for ionic diffusion [42] and strongly non-equilibrium phonon driving of ferroelectric materials, as we show below.
III Results
III.1 Dielectric Properties and Electrofreezing of Water
@@
Figure 2: Liquid water properties from ML-based molecular dynamics simulations under applied electric fields. (a) Calculated real and imaginary parts of the dielectric permittivity at 300 K in the 1-135000 GHz frequency range (see text). The gray points correspond to experimental data at 20C from Ref. [21]. Dashed lines serve as guides to the eye. (b) IR absorption spectra of liquid water at 300 K and varying electric-field strengths. (c) Average of the polarization along the direction of applied electric fields of varying strengths obtained from simulations with classical nuclei (red) and quantum nuclei (blue). Dashed lines represent fits assuming a model of classical non-interacting dipoles. (d) OH bond length and O-O nearest neighbor distance with varying field strengths in simulations with classical (red) and quantum (blue) nuclei. Also shown is the self-diffusion coefficient of water calculated from the root-mean-square displacements of oxygen atoms (classical nuclei) at varying field strengths. Shaded areas in all panels correspond to the standard error.
@@
We first apply the ML models we developed to study liquid water at room temperature and standard density. Water is of fundamental importance to life and, as such, it has been addressed numerous times in literature [49, 21, 50], with many peculiar features related to its dielectric properties and vibrational dynamics under applied electric fields still not fully resolved. The polarization of water results from a complex interplay between changes in the dipoles of individual water molecules and the correlated dynamics of multiple molecules in the liquid state. The results we present below were obtained by coupling the ML model we developed for the dipoles with a MLIP model for the energies and forces, see Section V.
@@
Our implementation allows nanosecond long simulations and the application of dynamic and static electric fields, going well beyond the time-scales reached by existing ab initio techniques. It is thus possible to efficiently simulate the frequency-dependent dielectric function $\varepsilon(\nu)$ of liquid water, as shown in Fig. 2a, in the range of $0.001-135$ THz ($0.03-4500$ cm-1). Spanning this large frequency range was made even more efficient through the combination of two simulation techniques that are allowed by our methodology. Above 100 GHz, the real and imaginary parts of $\varepsilon(\nu)$ were calculated through the fluctuation-dissipation relation [51, 52, 10] (see Supplementary Section S3), while the region below 100 GHz was obtained by a more efficient non-equilibrium simulation technique [51] (see Section V). We could then simulate the onset of the Debye relaxation, which reduces the value of $\varepsilon(\nu)$ at frequencies larger than 10 GHz [21]. The results shown in Fig. 2a agree quite well with experimental data [21] shown in the same figure, despite a slight underestimation of the static dielectric constant, which we attribute to the training of the $\tilde{\bm{\mu}}^{\text{MV}}$ model in this case. As we describe in Section V, the training was effectively based on mid-sized water clusters. We expect that training on larger periodic structures will solve this issue, better reproducing previous ab initio results [53]. Besides these differences, the results substantiate that our approach is able to qualitatively capture the atomistic mechanisms of the Debye relaxation in water and should be a useful tool to settle the debate about the spatial extent of dynamical molecular correlation involved in this process, without having to resort to empirical potentials for large system sizes [50, 53].
@@
In Fig. 2b we show the IR spectra of water under applied static electric fields, as obtained from the Fourier transform of dipole-autocorrelation time-series, simulated with molecular dynamics. The IR spectrum is closely related to ${\rm Im}\,\varepsilon(\nu)$ shown in Fig. 2a. The spectra in Fig. 2a are in excellent agreement with the ones recently reported in Ref. [17], which employed a similar ansatz of linear-coupling to the electric-field as we do here. With increasing field strength, there is a pronounced blue-shift of the libration band below $1000$ cm-1, a very small blue-shift of the bending band at $\approx 1700$ cm-1 and of the combination band at $\approx 2200$ cm-1, and a pronounced red-shift of the OH stretching band at $\approx 3600$ cm-1, effectively narrowing the spectral range. Such shifts are common features of the vibrational Stark effect in anharmonic systems [54]. The red-shift of the OH stretching band is due to a strengthening of H-bonds, evidenced by shorter O–O distances, as shown in Fig. 2d. The blue-shift of the libration band is due to the breaking of molecular rotational isotropy at increasing field strengths, as also discussed in Refs. [17, 23, 55]. Related to this blue-shift in the libration band, while the band we calculate is in excellent agreement with the one shown in Ref. [17] at all field strengths, both results do not contain the new peak just below $1000$ cm-1 at higher electric field strengths, reported by Cassone and coworkers and by Futera and English [23, 55], with ab initio MD simulations coupled to electric fields. This band is ascribed to a libration motion that is enhanced because of the preferential orientation of water along the electric field direction. As we also observe the preferential orientation of water molecules in our study (Fig. 2c), these differences must depend on the underlying methodology. We propose that the dynamics involved in the enhancement of this peak may be strongly dependent on the different underlying exchange-correlation functionals or on the inclusion of second-order electric-field coupling in the nuclear Hamiltonian, which are not explicitly considered in our work or in that of Ref. [17]. Because it is straightforward to augment our methodology to account for these terms, these considerations will be the subject of future work.
@@
We now analyze the impact of nuclear-quantum-effects (NQE) together with the application of electric fields in water. To the best of our knowledge, the importance of these effects at higher electric field strengths has not been previously discussed. In Fig. 2c we report the average polarization of liquid water along the field direction and at varying field strengths $\mathcal{E}$, from classical-nuclei and path-integral molecular dynamics (see Section V). While at zero-field there is no net polarization, the interaction with the field causes a net polarization to appear along the field direction. The increase of the average polarization for $\mathcal{E}\gtrsim 0.04$ V/Å  is non-linear, in agreement with Ref. [17]. This behavior can be understood by analyzing the thermodynamics of an ensemble of classical non-interacting dipoles in an external field. The analytical expression of the partition function of this system can be found in textbooks [56], and the total polarization of the system is expressed as a Langevin function that depends on $\mathcal{E}$ and the molecular dipoles. By fitting to this model (dashed lines in Fig. 2c), we conclude that NQEs make water molecules easier to polarize upon electric field application. The saturation value of the average polarization is higher when including NQE, because the enthalpic term of interaction with the field can better compensate the thermal entropy of the liquid.

### v2

@@ Abstract
The driving of vibrational motion by external electric fields is a topic of continued interest, due to the possibility of assessing new or metastable material phases with desirable properties. Here, we combine ab initio molecular dynamics within the electric-dipole approximation with machine-learning neural networks (NNs) to develop a general, efficient and accurate method to perform electric-field-driven nuclear dynamics for molecules, solids, and liquids. We train equivariant and autodifferentiable NNs for the interatomic potential and the dipole, modifying the model infrastructure to account for the multi-valued nature of the latter in periodic systems. We showcase the method by addressing property modifications induced by electric field interactions in a polar liquid and a polar solid from nanosecond-long molecular dynamics simulations with quantum-mechanical accuracy. For liquid water, we present a calculation of the dielectric function in the GHz to THz range and the electrofreezing transition, showing that nuclear quantum effects enhance this phenomenon. For the ferroelectric perovskite LiNbO3, we simulate the ferroelectric to paraelectric phase transition and the non-equilibrium dynamics of driven phonon modes related to the polarization switching mechanisms, showing that a full polarization switch is not achieved in the simulations.
@@ I Introduction
The interaction of matter with static and dynamical external electric fields plays a fundamental role in understanding the behavior of atoms, molecules, and complex materials. This interaction is exploited in fields as diverse as semiconductor nanotechnology [1] and enzymatic reactions [2]. In particular, there is a rising interest in tuning static and low-frequency dielectric fields to drive reversible phase transitions of materials [3, 4, 5, 6], with the goal of controlling mechanical properties, increasing the efficiency of ionic conductors and capacitors, or making better energy or memory storage devices [1, 2]. The large spatial extent and, depending on the phenomenon, the time-scale of many nanoseconds involved in the nuclear dynamics that define the response of a material to the application of such electric fields has been a long-standing challenge for quantum-mechanical first-principles atomistic simulations. While the physics of the matter-field coupling is known and appropriate simulation techniques already exist within the context of quantum-mechanical theories [7, 8, 9], simplified and non-transferable models [3] have often been used for describing these quantum phenomena at larger system sizes and longer time scales in complex systems, especially when targeting non-equilibrium dynamics.
@@
It is clear that the rapidly evolving techniques of machine-learning (ML) in atomistic simulations can be applied to this problem. ML models have already been proven to succeed in the prediction of energies, forces and electronic-structure properties of materials [10, 11] with an accuracy comparable to that of the underlying quantum electronic-structure method, but at a small fraction of the cost and often with a strongly reduced scaling with system size. A fundamental quantity to describe the interaction of matter with electric fields is the dipole and its derivative with respect to nuclear displacements. Several works have already proposed ML-based models for these quantities [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 16]. While these approaches are appropriate and successful in the investigated examples, i.e., close to structural equilibria, they can become inaccurate when large displacements take place, like in ionic diffusion or in non-equilibrium dynamics. As we will show in this work, an additional term arising from the multi-valued nature of the dipole in periodic systems sensitively influences the applicability of the ML model in these cases.
@@
By incorporating the multi-valued nature of the dipole in the ML model, we achieve a general framework that: (i) is generally applicable to different material classes (e.g. molecules, solids, liquids, and disordered systems) without relying on system-specific assumptions; (ii) keeps quantum-mechanical accuracy for electronic interactions and can include nuclear quantum effects; (iii) allows performing nanoseconds-long simulations including thousands of atoms in and out of equilibrium. We showcase the capabilities of our model by addressing liquid water and the solid ferroelectric perovskite LiNbO3 [23]. We calculate nanoseconds of first-principles-quality molecular dynamics coupled to static and time-dependent electric fields for these polar materials. For water, we show simulations of the dielectric function [24] spanning the GHz to the THz range and of the electrofreezing transition [25, 26, 27], finding an enhancement of this phenomenon due to nuclear quantum effects at larger field intensities. For LiNbO3, we show the calculation of the ferroelectric to paraelectric phase transition [28, 29, 30, 31] and of the full-dimensional non-equilibrium and non-linear driving of phonon-modes in this material. For this last example, incorporating the multi-valued nature of the dipole in the ML model is essential. It shows that full-dimensional simulations can better explain the experimental observation that an ultrafast polarization switch is only transiently and partially achieved [32, 33, 34].
@@
In the following, we will explain the ML model we develop, and discuss the rich phenomenology that these simulations can address together with the novel physical insights they can bring.
II Results
II.1 Machine-learning Model for Dipoles
@@
We start by characterizing the quantity we want to model. As further detailed in Section IV.2, the multi-valued nature of dipoles in periodic systems has the following consequences. If one imposes lattice periodicity on the dipole $\bm{\mu}$ by enforcing it to assume the same value at periodically-equivalent atomic sites, the function $\bm{\mu}(\bm{R})$ becomes discontinuous because this implies “switching branches” across the periodic boundaries, as shown in Fig. 1a. Instead, a smooth and continuous function is obtained by continuing to follow the same branch. Since the dipole then assumes different values for periodically-equivalent atomic positions, $\bm{\mu}(\bm{R})$ does not have periodic invariance with respect to the periodic boundary conditions.
@@
In other words, closed integrals do not vanish, but can take up (multiple) polarization quanta, depending on how often the periodic boundaries are crossed. We can quantify the change of the dipole $\Delta\bm{\mu}$ along a closed path $\gamma$ by a line integral. The result of this operation is non-trivial [35]:
@@
we can describe the more general case case where $N_{\rm at}$ atoms of the same oxidation state are displaced along the same direction. $\mathcal{N}_{I}$ is the difference between the values of $\mathcal{N}(s)$ upon crossing the periodic boundary. The values obtained for $\mathcal{N}_{I}$ are reported in Fig. 1b and Fig. 1c, and follow the expected oxidation numbers in these systems.
The previous considerations lead to practical consequences when building a ML model for $\bm{\mu}(\bm{R})$ of a periodic system. Currently, equivariant neural networks (ENNs) yield the best accuracy to machine-learn properties of materials [38]. Most commonly, ENNs take as inputs the tuple $\left(Z;\bm{R},\bm{a}_{\alpha}\right)$, where $Z$ is the atomic number, and model the target quantity as a function of the relative atomic positions $\Delta\bm{R}_{IJ}{{\color[rgb]{0,0,0}=\bm{R}_{I}-\bm{R}_{J}}}$, which are evaluated within the minimum image convention. This makes any target quantity a function of what we call the atomic environment $\bm{\mathcal{A}}$. Any target quantity which is a function of $\bm{\mathcal{A}}$ is invariant with respect to any displacement of atom to periodically equivalent sites.
Going back to Fig. 1a, because such models require that the function to be learned is differentiable, the discontinuous curve which is obtained by matching $\bm{\mu}$ to assume the same value at equivalent $\bm{\mathcal{A}}$ cannot be learned in its full domain. If one is only interested in small atomic displacements close to equilibrium, a usual ML model will provide accurate results for $\bm{\mu}$ in this restricted space. However, as one moves into $\bm{\mathcal{A}}(\bm{R})$ that are far from equilibrium, the ML model will start to change its slope, in order to obey $\bm{\mu}(\bm{\mathcal{A}}(\bm{R}+\bm{l}))=\bm{\mu}(\bm{\mathcal{A}}(\bm{R}))$. In an autodifferentiable model, this results in Born effective charges (BEC) $\bm{Z}^{*}$ (see Eqs. (8) in Section IV.1) and consequently forces that start to deviate from the correct value. If more data including environments far from equilibrium are added to the training set, the model will be unable to learn (see Supplementary Section S1 for a practical example).
@@
where $I$ runs over all atoms of the system and $\tilde{\bm{\mu}}$ can be handled by the usual learning procedure, as it depends only on $\bm{\mathcal{A}}$. From the definition of $\tilde{\bm{\mu}}^{\text{MV}}$ and $\bm{Z}^{*}$, we arrive at
@@
In the implementation we used in this work, we modified the output of the forward pass of MACE to include the second term on the right of Eq. (4). In this way, we always calculate $\tilde{\bm{\mu}}^{\text{MV}}$ in the training and use that also in the calculation of the loss-function. Training is successful and accurate even for datasets containing very out of equilibrium structures of H2O and LiNbO3, and $\bm{Z}^{*}$ obtained by autodifferentiation is also accurate (see Supplementary Section S1 and Section IV).
@@
The successes of previous ML models of dipoles and BEC are numerous [13, 17, 14, 41, 42, 21, 43, 19, 22, 12, 20]. They are successful because they either treat diffusive systems where the combined $\mathcal{N}$ (summing over all diffusing atoms) is zero, or address problems where the displacements along modes that result in significant variations of the dipole are not large and the system stays close to equilibrium along that coordinate. We note that models targeting to learn $\bm{Z}^{*}$ can also be problematic if they miss the last term in Eq. (5) and they are not guaranteed to obey the acoustic sum rule for this quantity [44]. The approach we present encompasses and extends these models to be generally applicable. This is for instance important for ionic diffusion [37] and strongly non-equilibrium phonon driving of ferroelectric materials, as we show below.
II.2 Dielectric Properties and Electrofreezing of Water
@@
Figure 2: Liquid water properties from ML-based molecular dynamics simulations under applied electric fields. (a) Calculated real and imaginary parts of the dielectric permittivity at 300 K in the 1-135000 GHz frequency range (see text). The gray points correspond to experimental data at 20C from Ref. [24]. Dashed lines serve as guides to the eye. (b) IR absorption spectra of liquid water at 300 K and varying electric-field strengths. (c) Average of the polarization along the direction of applied electric fields of varying strengths obtained from simulations with classical nuclei (red) and quantum nuclei (blue). Dashed lines represent fits assuming a model of classical non-interacting dipoles. (d) Top: OH bond length and O-O nearest neighbor distance with varying field strengths in simulations with classical (red) and quantum (blue) nuclei. Bottom: Self-diffusion coefficient of water calculated from the root-mean-square displacements of oxygen atoms at varying field strengths. Shaded areas in all panels correspond to the standard error.
@@
We first apply the ML models we developed to study liquid water at room temperature and standard density. Water is of fundamental importance to life and, as such, it has been addressed numerous times in literature [45, 24, 46], with many peculiar features related to its dielectric properties and vibrational dynamics under applied electric fields still not fully resolved. The polarization of water results from a complex interplay between changes in the dipoles of individual water molecules and the correlated dynamics of multiple molecules in the liquid state. The results we present below were obtained by coupling the ML model we developed for the dipoles with a machine-learning interatomic potential (MLIP) model for the energies and forces, see Section IV.
@@
Our implementation allows nanosecond long simulations and the application of dynamic and static electric fields, going well beyond the time-scales reached by existing ab initio techniques. It is thus possible to efficiently simulate the frequency-dependent dielectric function $\varepsilon(\nu)$ of liquid water, as shown in Fig. 2a, in the range of 0.0011350.001-135 THz (0.0345000.03-4500 cm-1). Spanning this large frequency range was made even more efficient through the combination of two simulation techniques that are allowed by our methodology. Above 100 GHz, the real and imaginary parts of $\varepsilon(\nu)$ were calculated through the fluctuation-dissipation relation [47, 48, 13] (see Supplementary Section S2), while the region below 100 GHz was obtained by a more efficient non-equilibrium simulation technique [47] (see Section IV). We could then simulate the onset of the Debye relaxation, which reduces the value of $\varepsilon(\nu)$ at frequencies larger than 10 GHz [24]. The results shown in Fig. 2a agree quite well with experimental data [24] shown in the same figure, despite a slight underestimation of the static dielectric constant, which we attribute to the training data included in the $\tilde{\bm{\mu}}^{\text{MV}}$ model in this case. As we show in Supplementary Section S3, adding dipoles computed in periodic bulk-water structures to the training dataset shifts the computed results closer to the experimental data and better reproduces previous ab initio results [49]. Besides these differences, the results substantiate that our approach is able to qualitatively capture the atomistic mechanisms of the Debye relaxation in water and should be a useful tool to settle the debate about the spatial extent of dynamical molecular correlation involved in this process, without having to resort to empirical potentials for large system sizes [46, 49].
In Fig. 2b we show the IR spectra of water under applied static electric fields, as obtained from the Fourier transform of dipole-autocorrelation time-series, simulated with molecular dynamics. The IR spectrum is closely related to ${\rm Im}\,\varepsilon(\nu)$ shown in Fig. 2a. The spectra in Fig. 2a are in excellent agreement with the ones recently reported in Ref. [20], which employed a similar ansatz of linear-coupling to the electric-field as we do here.
With increasing field strength, there is a pronounced blue-shift of the libration band below $1000$ cm-1, a very small blue-shift of the bending band at $\approx 1700$ cm-1 and of the combination band at $\approx 2200$ cm-1, and a pronounced red-shift of the OH stretching band at $\approx 3600$ cm-1, effectively narrowing the spectral range. The frequency shifts of these bands at the field strength of 0.15 V/Å  amount approximately to $113$, 77, and 75-75 cm-1, respectively. Such shifts are common features of the vibrational Stark effect in anharmonic systems [50]. The red-shift of the OH stretching band is due to a strengthening of H-bonds, evidenced by shorter O–O distances, as shown in Fig. 2d. The blue-shift of the libration band is due to the breaking of molecular rotational isotropy at increasing field strengths, as also discussed in Refs. [20, 26, 51].
@@
Related to this blue-shift in the libration band, while the band we calculate is in excellent agreement with the one shown in Ref. [20] at all field strengths, both results do not contain the new peak just below $1000$ cm-1 at higher electric field strengths, reported by Cassone and coworkers and by Futera and English [26, 51]. This band is ascribed to a libration motion that is enhanced because of the preferential orientation of water along the electric field direction. As we also observe the preferential orientation of water molecules in our study (Fig. 2c), these differences must depend on the underlying methodology. We propose that the dynamics involved in the enhancement of this peak may be strongly dependent on the different underlying exchange-correlation functionals or on the inclusion of second-order electric-field coupling in the nuclear Hamiltonian, which are not explicitly considered in our work or in that of Ref. [20]. Because it is straightforward to augment our methodology to account for these terms, these considerations will be the subject of future work.
@@
We now analyze the impact of nuclear quantum effects (NQE) together with the application of electric fields in water. To the best of our knowledge, the importance of these effects at higher electric field strengths has not been previously discussed. The narrowing of the IR spectrum with increasing field strength is more pronounced when accounting for NQE, as shown in Supplementary Section S4. The frequency shifts at 0.15 V/Å field strength amount approximately to $135$ cm-1 for the libration band and 175-175 cm-1 for the stretching band. It is interesting to look closer at the origin of this enhanced effect. In Fig. 2c we report the average polarization of liquid water along the field direction and at varying field strengths $\mathcal{E}$, from classical-nuclei and path integral molecular dynamics (PIMD). While at zero-field there is no net polarization, the interaction with the field causes a net polarization to appear along the field direction. The increase of the average polarization for $\mathcal{E}\gtrsim 0.04$ V/Å  is non-linear, in agreement with Ref. [20]. This behavior can be understood by analyzing the thermodynamics of an ensemble of classical non-interacting dipoles in an external field. The analytical expression of the partition function of this system can be found in textbooks [52], and the total polarization of the system is expressed as a Langevin function that depends on $\mathcal{E}$ and the molecular dipoles. By fitting to this model (dashed lines in Fig. 2c), we conclude that NQE make water molecules easier to polarize upon electric field application. The saturation value of the average polarization is higher when including NQE, because the enthalpic term of interaction with the field can better compensate the thermal entropy of the liquid.